OpenStudy (melissa_something):

Complete the square and write equation in standard form and give center and radius. -___-

OpenStudy (melissa_something):

x^2+y^2+2x+8y+17=19

OpenStudy (melissa_something):

I gathered like terms if anything x^2=2x, y^2+8y and 9-17=2

OpenStudy (anonymous):

2x (x+a)^2=x^2+2*x*a+a^2 since you have just 2x then a =1

OpenStudy (melissa_something):

Ok that part I got lost @AJ01

OpenStudy (anonymous):

okay you have Y^2 and 8y in the equation

OpenStudy (melissa_something):

Uh huh!

OpenStudy (anonymous):

stander form (y+b)^2

OpenStudy (anonymous):

(y+b)^2=y^2+2yb+b^2

OpenStudy (anonymous):

is that clear

OpenStudy (melissa_something):

um, I just don't know where you got b from, the equation right?

OpenStudy (melissa_something):

Teaching me math is really hard lol

OpenStudy (anonymous):

it is okay..... b: any number that i don't know yet

OpenStudy (anonymous):

ok

OpenStudy (melissa_something):

Ok. My bad trying to piece things together

OpenStudy (anonymous):

(y+b)^2=y^2+2yb+b^2 in the equation above you have 8y

OpenStudy (melissa_something):

What's the name of that formula?

OpenStudy (anonymous):

8y has to equal 2yb solve for b 2by=8y b=8y/2y=4

OpenStudy (melissa_something):

Why did you pick 8?

OpenStudy (anonymous):

Quadratic Formula Explained (x-a)^+(y-b)^=r^2

OpenStudy (anonymous):

(x-a)^2+(y-b)^2=r^2 x^2-2ax+a^2+y^2-2by+b^2=r^2

OpenStudy (anonymous):

the equation you have like this one i want put back to the stander form..... so 2by has to equal something with y

OpenStudy (melissa_something):

Woah

OpenStudy (melissa_something):

Okay!

OpenStudy (anonymous):

x^2+y^2+2x+8y+17=19 you need (1) for x form and (16) for the y form (x+1)^2+(y+4)^2=19

OpenStudy (melissa_something):

OMG MY BAD the 19 is a 9!

OpenStudy (anonymous):

okay then 3^2

OpenStudy (anonymous):

(x+1)^2+(y+4)^2=3^2 Explained this X^2+2x+1+y^2+8y+16=9 x62+2x+y^2+8y+17=9

OpenStudy (melissa_something):

That is all?